Use the dynamic programming technique to find an optimal parenthesization of a matrix-chain product whose sequence of dimensions is <5, 8, 40, 10, 20, 6>.

Matrix Dimension

A1 5 * 8

A2 8*40

A3 40*10

A4 10*20

A5 20*6
Answer #1

[A1(A2A3)(A4A5)]=5100

explanation

A2*A3=8*40*10=3200

A1(A2*A3)=5*8*10=400

(A4*A5)=10*20*6=1200

[A1(A2*A3)(A4*A5)]=5*10*6=300

therefore total=3200+400+1200+300